OpenStudy (anonymous):

how to convert 0-5 range to 70-130 range

OpenStudy (anonymous):

IS THAT THE FULL QUESTION?

OpenStudy (anonymous):

ya...0 must get converted to 70 and 5 to 130...and the ranges in between must be converted accordingly..

OpenStudy (unklerhaukus):

multiply by 12 and add 70

OpenStudy (anonymous):

12???

OpenStudy (unklerhaukus):

0...5 times 12 plus 70 0...60 plus 70 70...130

OpenStudy (unklerhaukus):

the range of 70-130 is 60 60/5 = 12

OpenStudy (anonymous):

actually its a conversion program reg our project..so v need to convert 0-5 volts to a range of 70-130mm of Hg

OpenStudy (unklerhaukus):

\[V(v)=(12v+70)[\text{mmHg}]\]

OpenStudy (anonymous):

oh..got it thanx

OpenStudy (unklerhaukus):

\[V(0)=(12(0)+70)[\text{mmHg}]\]\[\qquad=(0+70)[\text{mmHg}]\]\[\quad=70[\text{mmHg}]\] \[V(5)=(12(5)+70)[\text{mmHg}]\]\[\qquad=(60+70)[\text{mmHg}]\]\[\quad=130[\text{mmHg}]\]

OpenStudy (mayankdevnani):

@UnkleRhaukus good work......

OpenStudy (unklerhaukus):

\[\color{grey}{\ddot\smile}\]